If the speed sound in air is 330m/s at 0°C. What is the speed at 15°C and 100°C


1
Expert's answer
2021-03-09T15:27:48-0500

We can find the speed of sound from the formula:


"v=\sqrt{\dfrac{\gamma RT}{M}}."

Let's calculate the speed of sound at 15°C:


"v=\sqrt{\dfrac{1.4\cdot8.314\ \dfrac{J}{K\cdot mol}\cdot288.15\ K}{0.02896\ \dfrac{kg}{mol}}}=340\ \dfrac{m}{s}."

Let's calculate the speed of sound at 100°C:


"v=\sqrt{\dfrac{1.4\cdot8.314\ \dfrac{J}{K\cdot mol}\cdot373.15\ K}{0.02896\ \dfrac{kg}{mol}}}=387.3\ \dfrac{m}{s}."